# Learning to Steer Markovian Agents under Model Uncertainty

### v1

@@ Abstract
Designing incentives for an adapting population is a ubiquitous problem in a wide array of economic applications and beyond. In this work, we study how to design additional rewards to steer multi-agent systems towards desired policies without prior knowledge of the agents’ underlying learning dynamics. We introduce a model-based non-episodic Reinforcement Learning (RL) formulation for our steering problem. Importantly, we focus on learning a history-dependent steering strategy to handle the inherent model uncertainty about the agents’ learning dynamics. We introduce a novel objective function to encode the desiderata of achieving a good steering outcome with reasonable cost. Theoretically, we identify conditions for the existence of steering strategies to guide agents to the desired policies. Complementing our theoretical contributions, we provide empirical algorithms to approximately solve our objective, which effectively tackles the challenge in learning history-dependent strategies. We demonstrate the efficacy of our algorithms through empirical evaluations [^1] The code of all the experiments in this paper can be found in https://github.com/jiaweihhuang/Steering_Markovian_Agents..
@@ 1 Introduction
This paper focuses on situations when an external “mediator” exists, who can influence and steer the agents’ learning dynamics by modifying the original rewards via additional incentives. Following the works of Zhang et al., (2023) and Canyakmaz et al., (2024) (see more discussion in Sec. 1.1), this kind of mediator can be conceptualized in various ways. In particular, we can think of a social planner who provides monetary incentives for joint ventures or for adoption of an innovative technology via individual financial subsidies. As illustrated on the RHS of Fig. 1-(b), with suitable steering, agents’ dynamics can be directed to the best outcome. In this paper, we study the problem from the mediator perspective, and focus on developing steering strategies for Markovian agents whose policy learning dynamics only depend on their current policy and on their (modified) reward function. Our primary objective is to steer the agents to some desired policies, that is, to minimize the steering gap vis-a-vis the target outcome. As a secondary objective, the payments to agents regarding the steering rewards should be reasonable, that is, the steering cost should be low.
@@
How can we learn desired steering strategies for Markovian agents
@@
We consider a model-based setting where the mediator can get access to a model class $\mathcal{F}$ containing the agents’ true learning dynamics $f^{*}$. We summarize our main results in Table 1 and briefly highlight our key contributions as follow:
Table 1: A Summary of Main Results
A New Objective Function in Obj. (1)
Settings Existence (Sec. 4) Algorithms (Sec. 5)
=$\mathcal{F}=\{f^{*}\}$ (Special case: known $f^{*}$) Thm. 4.2 Any RL/Control Algorithm
$|\mathcal{F}|$ is small Prop. 4.5 Alg. 4 (A model belief-state method)
$|\mathcal{F}|$ is large Proc. 2 (The FETE framework)
@@ 1.1 Closely Related Works
We discuss the works most closely related to ours in this section, and defer the others to Appx. C.2.
Steering Learning Dynamics The ‘steering problem’ as we study it in this paper was first introduced by Zhang et al., (2023). They consider the case of no-regret learners who may possibly know the mediator’s steering strategy and can be arbitrarily adversarial. Their focus is on steering agents such that the average policy converges to the target NE while the accumulative budget is sublinear. When the desired NE is not pure, Zhang et al., (2023) further require the mediator to be able to “give advice” to the players. By contrast, we focus on a broader class of Markovian agents with a finite-horizon steering setup, with focus on the steering gap of the terminal policy and the cumulative steering cost. Besides, our mediator can steer only by modifying the agents’ reward functions.
Perhaps the closest to ours is a concurrent work by Canyakmaz et al., (2024). They experimentally investigate the use of control methods to direct game dynamics towards desired outcomes, in particular allowing for model uncertainty. Our work complements their empirical results in three main ways. First, we handle the model uncertainty in a somewhat more principled way by proposing a concrete non-episodic RL formulation for the steering problem and a suitable learning objective (Obj. (1)), where we explicitly learn a history-dependent steering strategy. Second, we develop novel theory regarding the existence of strategies with low steering gap. Third, on the algorithmic level, Canyakmaz et al., (2024) consider a two-phase (exploration + exploitation) framework called SIAR-MPC which is similar to our FETE framework for large model sets. However, they employ random noise-based exploration, while we consider a more advanced exploration strategy (Sec. 5.2) that results in significantly higher exploration efficiency in experiments (Sec. 6). Besides, when the model set is small, we contribute a belief-state based algorithm that can exactly solve our Obj. (1). As suggested by Prop. 3.3, it is superior to their two-phase framework in terms of the steering cost while achieving a low steering gap.
Episodic RL and Non-Episodic RL
@@ 2 Preliminary
In the following, we formally define the finite-horizon Markov Game that we will focus on. We summarize all the frequently used notations in this paper in Appx. B.
@@ 3 The Problem Formulation of the Steering Markovian Agents
We first introduce our definition of Markovian agents. Informally, the policy updates of Markovian agents are independent of the interaction history conditioning on their current policy and observed rewards. This subsumes a broader class of popular policy-based methods as concrete examples (Giannou et al.,, 2022; Ding et al.,, 2022; Xiao,, 2022; Daskalakis et al.,, 2020).
@@
Here $\psi:=\{\psi_{t}\}_{t\in[T]}$ denotes the mediator’s “steering strategy” to generate $u_{t}$. We consider history-dependent strategies to handle the model uncertainty, which we will explain later. Besides, ${\bm{u}}_{t}:=\{u_{t,h}^{n}\}_{h\in[H],n\in[N]}$, where $u^{n}_{t,h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,U_{\max}]$ is the steering reward for agent $n$ at game horizon $h$ and steering step $t$. $U_{\max}<+\infty$ denotes the upper bound for the steering reward. For practical concerns, we follow Zhang et al., (2023) and constrain the steering reward to be non-negative.
@@
Steering Dynamics as a Markov Decision Process (MDP) Given a game $G$, the agents’ dynamics $f$ and $(\eta^{\text{cost}},\eta^{\text{goal}})$, the steering dynamics can be modeled by a finite-horizon MDP. $M:=\{{\bm{\pi}}_{1},T,\Pi,\mathcal{U},f,(\eta^{\text{cost}},\eta^{\text{goal}})\}$ with initial state ${\bm{\pi}}_{1}$, horizon length $T$, state space $\Pi$, action space $\mathcal{U}:=[0,U_{\max}]^{HN|\mathcal{S}||\mathcal{A}|}$, stationary transition $f$, running reward $\eta^{\text{cost}}$ and terminal reward $\eta^{\text{goal}}$. For completeness, we defer to Appx. C.3 for an introduction of finite-horizon MDP
@@ 4.1 Existence when $f^{*}$ is Known: Natural Policy Gradient as an Example
In this section, we focus on a popular choice of learning dynamics called Natural Policy Gradient (NPG) dynamics (Kakade,, 2001; Agarwal et al.,, 2021) (a.k.a. the replicator dynamics (Schuster and Sigmund,, 1983)) with direct policy parameterization. NPG is a special case of the Policy Mirror Descent (PMD) (Xiao,, 2022). For the readability, we stick to NPG in the main text, and in Appx. E.1, we formalize PMD and extend the results to the general PMD, which subsumes other learning dynamics, like the online gradient ascent (Zinkevich,, 2003).
@@
Our result is strong in indicating the existence of a steering path for any feasible initialization. The proof is based on construction. The basic idea is to design the ${\bm{u}}_{t}$ so that $A^{{\bm{\pi}}_{t}}_{|{\bm{r}}+{\bm{u}}_{t}}\propto\log\frac{{\bm{\pi}}^{*}}{{\bm{\pi}}_{t}}$, for some target policy ${\bm{\pi}}^{*}\in\Pi^{+}$ (approximately) maximizing $\eta^{\text{goal}}$, then we can guarantee the convergence of ${\bm{\pi}}_{t}$ towards ${\bm{\pi}}^{*}$ under Assump. B. The main challenge here would be the design of $u_{t}$. We defer the details and the formal statements to Appx. E.
@@ 4.2 Existence when $f^{*}$ is Unknown: the Identifiable Model Class
[A Sufficient Condition for Existence] Given any $\varepsilon>0$, $\Psi^{\varepsilon}_{T}(\mathcal{F};{\bm{\pi}}_{1})\neq\emptyset$ [^5] Here we highlight the dependence on initial policy, model, and time for clarity (see Footnote 4), if $\exists~{}\widetilde{T}<T$, s.t., (1) $\mathcal{F}$ is $(\frac{\varepsilon}{2\eta_{\max}},\widetilde{T})$-identifiable, (2) $\Psi^{\varepsilon/2}_{T-\widetilde{T}}(\mathcal{F};{\bm{\pi}}_{\widetilde{T}})\neq\emptyset$ for any possible ${\bm{\pi}}_{\widetilde{T}}$ generated at step $\widetilde{T}$ during the steering.
We conclude this section by noting that, by Thm. 4.2, the above condition (2) is realistic for NPG (or more general PMD) dynamics. The proofs for all results in this section are deferred to Appx. F.
@@ 5 Learning (Approximately) Optimal Steering Strategy
As a special case, when the model is known, i.e. =$\mathcal{F}=\{f^{*}\}$, by the Markovian property, Obj. (1) reduces to a normal RL objective, and a state-dependent steering strategy $\psi:\Pi\rightarrow\mathcal{U}$ is already enough. For completeness, we include the algorithm but defer to Alg. 3 in Appx. C.4. In the rest of this section, we focus on the general case $|\mathcal{F}|>1$. In Sec. 5.1, we investigate the solutions when $|\mathcal{F}|$ is small, and in Sec. 5.2, we study the more challenging case when $|\mathcal{F}|$ is large.
@@ 5.1 Small Model Class: Dynamic Programming with Model Belief State
A Partially Observable MDP Perspective In fact, we can interpret Obj. (1) as learning the optimal policy in a POMDP, in which the hidden state is $({\bm{\pi}}_{t},f)$, i.e. a tuple containing the policy and the hidden model $f$ uniformly sampled from $\mathcal{F}$, and the mediator can only partially observe the policy ${\bm{\pi}}_{t}$. It is well-known that any POMDP can be lifted to the belief MDP, where the state is the belief state of the original POMDP. Then, the optimal policy in the belief MDP is exactly the optimal history-dependent policy in the original POMDP (Ibe,, 2013). In our case, for each step $t\in[T]$, the belief state is $({\bm{\pi}}_{t},b_{t})$, where $b_{t}:=[\Pr(f|\{{\bm{\pi}}_{t^{\prime}},{\bm{u}}_{t^{\prime}}\}_{t^{\prime}=1}^{t},{\bm{\pi}}_{t})]_{f\in\mathcal{F}}$ is the “model belief state” defined to be the posterior distribution of models given the history of observations and actions. When $|\mathcal{F}|$ is small, the model belief state $b_{t}\in\mathcal{R}^{|\mathcal{F}|}$ is low dimensional and computable. Learning $\psi^{*}$ is tractable by running any RL algorithm on the lifted MDP. In Proc. 1, we show how to steer in this setting. We defer the detailed algorithm of learning such belief-state dependent strategy to Alg. 4 in Appx. C.5.
@@ 3 Learn an exploration strategy
Deploy $\psi^{\text{Explore}}$ to steer the real agents and collect $\{{\bm{\pi}}_{1},{\bm{u}}_{1},...,{\bm{\pi}}_{\widetilde{T}},{\bm{u}}_{\widetilde{T}},{\bm{\pi}}_{\widetilde{T}+1}\}$

### v2

@@ Abstract
Designing incentives for an adapting population is a ubiquitous problem in a wide array of economic applications and beyond. In this work, we study how to design additional rewards to steer multi-agent systems towards desired policies without prior knowledge of the agents’ underlying learning dynamics. Motivated by the limitation of existing works, we consider a new and general category of learning dynamics called Markovian agents. We introduce a model-based non-episodic Reinforcement Learning (RL) formulation for our steering problem. Importantly, we focus on learning a history-dependent steering strategy to handle the inherent model uncertainty about the agents’ learning dynamics. We introduce a novel objective function to encode the desiderata of achieving a good steering outcome with reasonable cost. Theoretically, we identify conditions for the existence of steering strategies to guide agents to the desired policies. Complementing our theoretical contributions, we provide empirical algorithms to approximately solve our objective, which effectively tackles the challenge in learning history-dependent strategies. We demonstrate the efficacy of our algorithms through empirical evaluations [^1] The code of all the experiments in this paper can be found in https://github.com/jiaweihhuang/Steering_Markovian_Agents..
@@ 1 Introduction
This paper focuses on situations when an external “mediator” exists, who can influence and steer the agents’ learning dynamics by modifying the original rewards via additional incentives. This kind of mediator can be conceptualized in various ways. In particular, we can think of a social planner who provides monetary incentives for joint ventures or for adoption of an innovative technology via individual financial subsidies. As illustrated on the RHS of Fig. 1-(b), with suitable steering, agents’ dynamics can be directed to the best outcome. Our primary objective is to steer the agents to some desired policies, that is, to minimize the steering gap vis-a-vis the target outcome. As a secondary objective, the payments to agents regarding the steering rewards should be reasonable, that is, the steering cost should be low.
@@
To our knowledge, Zhang et al., (2023) is the first work studying a similar steering problem as ours. They assume the agents are no-regret learners and may act in arbitrarily adversarial ways. In some natural settings, such assumption may not be practical, because no-regret criterion typically requires careful processing of the entire history of learning. In settings with limited cognitive resources and bounded rationality, it is natural to favor models where the agents only process a subset of the available information (Camerer,, 2011). In particular, humans have been widely shown to rely overproportionally on recent experiences in decision making, known as ‘recency bias’ (Costabile and Klein,, 2005; Page and Page,, 2010; Durand et al.,, 2021). Besides, there is evidence that behavioral dynamics that only rely on the most recent experience are able to fit behavioral data well in certain situations (Mäs and Nax,, 2016). Motivated by these insights, we therefore study steering a different category of learning dynamics called Markovian agents, where the agents’ policy updates only depend on their current policy and the (modified) reward function. Our model complements the prior work on no-regret agents, and serves as the first abstraction of behavior based on limited cognitive abilities with recency bias in steering setting. Theoretically, Markovian agents subsumes a broader class of popular policy-based methods as concrete examples (Giannou et al.,, 2022; Ding et al.,, 2022), which are not covered by no-regret assumptions. We also note that a concurrent work (Canyakmaz et al.,, 2024) considers a similar setting as ours, and we defer to Sec. 1.1 for further discussion.
@@
How to learn desired steering strategies for Markovian agents
@@
We consider a model-based setting where the mediator can get access to a model class $\mathcal{F}$ containing the agents’ true learning dynamics $f^{*}$. We summarize and highlight our key contributions as follow:
@@ 1.1 Closely Related Works
We discuss the works most closely related to ours in this section, and defer the others to Appx. B.2.
Steering Learning Dynamics As mentioned in the introduction, Zhang et al., (2023) are the first to introduce the “steering problem”, but their setting differs quite fundamentally from ours in several key aspects. Firstly, they assume that agents behave as no-regret and arbitrarily adversarial learners, which may be unrealistic in settings with limited information and feedback, and owed to agents’ limited cognitive resources (Camerer,, 2011) including recency bias (Costabile and Klein,, 2005). Motivated by this, we instead focus on a broad class of Markovian dynamics. Secondly, the mediator’s objective in Zhang et al., (2023) is to steer agents such that the average policy converges to the target NE while maintaining a sublinear accumulative budget, motivated by their infinite-horizon setup. In contrast, we consider the finite-horizon setting, and therefore, we are concerned with minimizing the steering gap of the terminal policy and the cumulative steering cost. Thirdly, when the desired NE is not pure, Zhang et al., (2023) require the mediator to be able to “give advice” to the players to facilitate coordination, while we do not allow the mediator to do this. Because of these differences, the methods and results obtained by Zhang et al., (2023) and us are not directly comparable, yet complement one another depending on application considered.
Perhaps the closest to ours is a concurrent work by Canyakmaz et al., (2024), which contributes empirical investigation on the use of control methods to direct game dynamics towards desired outcomes, in particular allowing for model uncertainty. Although not formally specified in Canyakmaz et al., (2024), they consider a similar finite-horizon non-episodic setup as ours. However, they only consider history-independent steering strategy, which can result in sub-optimal performance in this finite-horizon setup, because one-step observations may not provide sufficient information to tackle model uncertainty. As our main contribution compared with their work, we point out that, one should, in principle, employ history-dependent steering strategies, since history can serve as sufficient information set for decision making under uncertainty. This leads to significant differences in the design principles of our algorithms compared with Canyakmaz et al., (2024). Concretely, we propose a learning objective for history-dependent strategies in Obj. (1), and two algorithms for low uncertainty (small $\mathcal{F}$) and high uncertainty (large $\mathcal{F}$) settings, respectively. In the former case, we contribute a belief-state based algorithm that can exactly solve Obj. (1), offering a stronger solution than Canyakmaz et al., (2024) due to the theoretical guarantee in Prop. 3.3. For the latter, although both our FETE and SIAR-MPC (Canyakmaz et al.,, 2024) share a two-phase (exploration + exploitation) structure, ours represents a more general framework with a more advanced exploration strategy (see more explanation in Sec. 5.2). Besides, we develop additional novel theory regarding the existence of strategies with low steering gap.
@@ 2 Preliminary
In the following, we formally define the finite-horizon Markov Game that we will focus on. We summarize all the frequently used notations in this paper in Appx. A.
@@ 3 The Problem Formulation of the Steering Markovian Agents
We first introduce our definition of Markovian agents. Informally, the policy updates of Markovian agents are independent of the interaction history conditioning on their current policy and observed rewards. This encompass a broader class of popular policy-based methods as concrete examples (Giannou et al.,, 2022; Ding et al.,, 2022; Xiao,, 2022; Daskalakis et al.,, 2020).
@@
Here $\psi:=\{\psi_{t}\}_{t\in[T]}$ denotes the mediator’s “steering strategy” to generate $u_{t}$. We consider history-dependent strategies to handle the model uncertainty, which we will explain later. Besides, ${\bm{u}}_{t}:=\{u_{t,h}^{n}\}_{h\in[H],n\in[N]}$, where $u^{n}_{t,h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,U_{\max}]$ is the steering reward for agent $n$ at game horizon $h$ and steering step $t$. $U_{\max}<+\infty$ denotes the upper bound for the steering reward. For practical concerns, we follow the standard constraints that the steering rewards are non-negative.
@@
Steering Dynamics as a Markov Decision Process (MDP) Given a game $G$, the agents’ dynamics $f$ and $(\eta^{\text{cost}},\eta^{\text{goal}})$, the steering dynamics can be modeled by a finite-horizon MDP. $M:=\{{\bm{\pi}}_{1},T,\Pi,\mathcal{U},f,(\eta^{\text{cost}},\eta^{\text{goal}})\}$ with initial state ${\bm{\pi}}_{1}$, horizon length $T$, state space $\Pi$, action space $\mathcal{U}:=[0,U_{\max}]^{HN|\mathcal{S}||\mathcal{A}|}$, stationary transition $f$, running reward $\eta^{\text{cost}}$ and terminal reward $\eta^{\text{goal}}$. For completeness, we defer to Appx. B.3 for an introduction of finite-horizon MDP
@@ 4.1 Existence when $f^{*}$ is Known: Natural Policy Gradient as an Example
In this section, we focus on a popular choice of learning dynamics called Natural Policy Gradient (NPG) dynamics (Kakade,, 2001; Agarwal et al.,, 2021) (a.k.a. the replicator dynamics (Schuster and Sigmund,, 1983)) with direct policy parameterization. NPG is a special case of the Policy Mirror Descent (PMD) (Xiao,, 2022). For the readability, we stick to NPG in the main text, and in Appx. D.1, we formalize PMD and extend the results to the general PMD, which subsumes other learning dynamics, like the online gradient ascent (Zinkevich,, 2003).
@@
Our result is strong in indicating the existence of a steering path for any feasible initialization. The proof is based on construction. The basic idea is to design the ${\bm{u}}_{t}$ so that $A^{{\bm{\pi}}_{t}}_{|{\bm{r}}+{\bm{u}}_{t}}\propto\log\frac{{\bm{\pi}}^{*}}{{\bm{\pi}}_{t}}$, for some target policy ${\bm{\pi}}^{*}\in\Pi^{+}$ (approximately) maximizing $\eta^{\text{goal}}$, then we can guarantee the convergence of ${\bm{\pi}}_{t}$ towards ${\bm{\pi}}^{*}$ under Assump. B. The main challenge here would be the design of $u_{t}$. We defer the details and the formal statements to Appx. D.
@@ 4.2 Existence when $f^{*}$ is Unknown: the Identifiable Model Class
[A Sufficient Condition for Existence] Given any $\varepsilon>0$, $\Psi^{\varepsilon}_{T}(\mathcal{F};{\bm{\pi}}_{1})\neq\emptyset$ [^4] Here we highlight the dependence on initial policy, model, and time for clarity (see Footnote 3), if $\exists~{}\widetilde{T}<T$, s.t., (1) $\mathcal{F}$ is $(\frac{\varepsilon}{2\eta_{\max}},\widetilde{T})$-identifiable, (2) $\Psi^{\varepsilon/2}_{T-\widetilde{T}}(\mathcal{F};{\bm{\pi}}_{\widetilde{T}})\neq\emptyset$ for any possible ${\bm{\pi}}_{\widetilde{T}}$ generated at step $\widetilde{T}$ during the steering.
We conclude this section by noting that, by Thm. 4.2, the above condition (2) is realistic for NPG (or more general PMD) dynamics. The proofs for all results in this section are deferred to Appx. E.
@@ 5 Learning (Approximately) Optimal Steering Strategy
As a special case, when the model is known, i.e. =$\mathcal{F}=\{f^{*}\}$, by the Markovian property, Obj. (1) reduces to a normal RL objective, and a state-dependent steering strategy $\psi:\Pi\rightarrow\mathcal{U}$ is already enough. For completeness, we include the algorithm but defer to Alg. 3 in Appx. B.4. In the rest of this section, we focus on the general case $|\mathcal{F}|>1$. In Sec. 5.1, we investigate the solutions when $|\mathcal{F}|$ is small, and in Sec. 5.2, we study the more challenging case when $|\mathcal{F}|$ is large.
@@ 5.1 Small Model Class: Dynamic Programming with Model Belief State
A Partially Observable MDP Perspective In fact, we can interpret Obj. (1) as learning the optimal policy in a POMDP, in which the hidden state is $({\bm{\pi}}_{t},f)$, i.e. a tuple containing the policy and the hidden model $f$ uniformly sampled from $\mathcal{F}$, and the mediator can only partially observe the policy ${\bm{\pi}}_{t}$. It is well-known that any POMDP can be lifted to the belief MDP, where the state is the belief state of the original POMDP. Then, the optimal policy in the belief MDP is exactly the optimal history-dependent policy in the original POMDP (Ibe,, 2013). In our case, for each step $t\in[T]$, the belief state is $({\bm{\pi}}_{t},b_{t})$, where $b_{t}:=[\Pr(f|\{{\bm{\pi}}_{t^{\prime}},{\bm{u}}_{t^{\prime}}\}_{t^{\prime}=1}^{t},{\bm{\pi}}_{t})]_{f\in\mathcal{F}}$ is the “model belief state” defined to be the posterior distribution of models given the history of observations and actions. When $|\mathcal{F}|$ is small, the model belief state $b_{t}\in\mathcal{R}^{|\mathcal{F}|}$ is low dimensional and computable. Learning $\psi^{*}$ is tractable by running any RL algorithm on the lifted MDP. In Proc. 1, we show how to steer in this setting. We defer the detailed algorithm of learning such belief-state dependent strategy to Alg. 4 in Appx. B.5.
@@ 3 Learn an exploration strategy
4 Deploy $\psi^{\text{Explore}}$ to steer the real agents and collect $\{{\bm{\pi}}_{1},{\bm{u}}_{1},...,{\bm{\pi}}_{\widetilde{T}},{\bm{u}}_{\widetilde{T}},{\bm{\pi}}_{\widetilde{T}+1}\}$
@@
Comparison with Canyakmaz et al., (2024) Although both SIAR-MPC in Canyakmaz et al., (2024) and our FETE (Procedure 2) adopt a first-explore-then-exploit structure, FETE is more general and more effective. Both algorithms have three main components: exploration strategy, model estimation strategy and exploitation strategy, and we inspect our advantages from these three aspects. (i) Exploration strategy: SIAR-MPC uses noise-based random exploration, whereas we adopt a more strategic approach, which uses the identification success rate as a signal to learn the exploration policy. Empirical results in Sec. 6.2 demonstrate the higher efficiency of our methods. (ii) Model estimation strategy: SIAR-MPC estimates the hidden model by solving a regression problem with constraints (Eq. (8) in Canyakmaz et al., (2024)), while we solve a MLE objective. In fact, our MLE objective is more general and can recover the regression problem in SIAR-MPC, if we consider a model class $\mathcal{F}$ that includes Gaussian noise perturbed dynamics with the side-information constraints introduced in Canyakmaz et al., (2024). (iii) Exploitation strategy: As a general framework, our FETE does not restrict how to compute the exploitation strategy $\psi^{Exploit}$. As we suggest in paper, any RL or control method can be used, including the MPC approach in Canyakmaz et al., (2024).